OpenStudy (anonymous):

Find the height at which g is reduced to g/2.

OpenStudy (anonymous):

Let g/2 = k. mg = GmM/R^2 => g = GM/R^2. g/k = 2. So , g/k = 2 = H^2/R^2 , H = 1.414*R = 9049.6 Km. Where R is the radius of earth and H is the height at which g becomes g/2.